OpenStudy (anonymous):

Michelle draws a card from a standard deck of 52 cards. She replaces the card and draws a second card. What is the probability that she draws a red card and then a black card?

OpenStudy (anonymous):

find each probability and them multiply them together

OpenStudy (anonymous):

P(red) = 26/52 = 1/2 etc as satellite says